40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1626     Entry time: Thu Apr 7 21:52:43 2016
Author: Antonio and Tara 
Type: DailyProgress 
Category: PMC 
Subject: PMC number 2 

Today we have tested the second PMC that we have assembled.

 

We have obtained the 00 mode;

We measured the finesse (~300);

We have scanned the PMC PZT and we see 3 FSR;

 

We also noted again the presence of a second 00 mode and we confirmed that it was due to laser mode hopping. This has been checked

by injecting a voltage (0.2V) in the temperature input of the laser. Once the voltage was applied the second mode disappeard.

 

We think that it is enough PMC testing for now. I am tempted to assemble a third PMC, while we should try to fix the glass PMC.

We need to work on the lab organization too.

ELOG V3.1.3-